OpenStudy (anonymous):

Let f(x) = 27x5 - 33x4 - 21x3 and g(x) = 3x2

OpenStudy (anonymous):

\[\frac{ f(x) }{g(x)}\]

OpenStudy (anonymous):

find that

OpenStudy (anonymous):

same thing, just factor out a 3x^2

OpenStudy (anonymous):

9x^3

OpenStudy (anonymous):

- 11x^2 - 7x

OpenStudy (anonymous):

f(x) can be rewritten as\[f(x) = 3x^2(9x^3 - 11x^2 - 7x)\] so dividing that by 3x^2 will be just what we factored out

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

and what do we do with \frac{ f(x) }{g(x)}

OpenStudy (anonymous):

\[\frac{ f(x) }{ g(x) }\]

OpenStudy (anonymous):

ohh! Thats just telling us to divide!

OpenStudy (anonymous):

f(x)/g(x) = 9x^3 - 11x^2 - 7x

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

Thank you!